# Multimodal AI predicts clinical outcomes of drug combinations from preclinical data

### v1

@@ Discussion
Combination therapies are essential for treating complex diseases such as hypertension, cancer, and infectious diseases, but current models often fail to translate preclinical data to predict clinical drug combination outcomes due to their focus on limited types of preclinical data, such as structural or target information. These approaches struggle to capture biological interactions that arise from multi-drug regimens. Integrating diverse data modalities, including structural data, pathways, cell viability, and transcriptomics, addresses this gap, offering a more holistic approach to predicting drug interactions. We developed Madrigal, a universal multimodal model that predicts the effects of drug combinations in 953 clinical outcomes and 21,842 compounds, including approved drugs and those in development. Madrigal surpasses existing single-modality models and state-of-the-art techniques to predict adverse drug interactions. Madrigal predicts clinical transporter-mediated DDIs and safety of PARPi anti-cancer combinations and chronic conditions such as T2D and MASH. Madrigal also extends to personalized cancer therapies by predicting individualized drug combination outcomes from BeatAML genomic profiles and patient-derived xenograft models.
@@
Madrigal can be used for preclinical predictions. When predicting adverse event risks of drug combinations to be tested in the ComboMATCH trial [60], Madrigal predicts elevated risk for eight combinations (Supplementary Fig. S8c). We observe correlations between proteomic changes following drug perturbations[116] and drug-drug interactions, reflecting broader pathway effects (Supplementary Fig. S8a, $p$-value $<$ 0.001 between neighboring groups, Mann-Whitney U). Proteomic changes can partially explain the similarities between drug embeddings generated by Madrigal, even when controlling for similarities in the target profile, suggesting that Madrigal captures off-target and pathway effects (Supplementary Fig. S8b).
Multimodal models such as Madrigal can help select candidates for combination testing and offer new insights into the biological mechanisms underlying drug synergies. Flexibility in modeling outcomes and encoding drugs is crucial for the broader applicability of safety screening pipelines. By integrating Madrigal with an LLM, users can describe and predict clinical effects in natural language, including those not captured by existing controlled vocabularies or standardized terminologies. This flexibility helps mitigate some limitations in drug combination effects data, such as incomplete annotations for specific outcomes (e.g., neutropenia for (olaparib + paclitaxel)) that, while reported in patients [117, 118], are not yet captured in drug effect training datasets. Data biases arising from clinical practice can influence predictions, for example, when specific drugs, such as SGLT2 inhibitors, are preferentially prescribed to specific populations, such as type II diabetic patients with heart failure [119]. This could have led to different groups of drug combinations being taken by different populations with distinct underlying risks of certain adverse events. The ability to input flexible, free-form descriptions in Madrigal-LLM could help address these gaps by enabling scientists to benchmark candidate drug combinations, potentially increasing both the success rate in drug development and the number of patients who tolerate these combinations.
One limitation of Madrigal lies in the indication-agnostic nature of its training data and predictions. Although the model is trained on datasets with coverage of drug combinations and outcomes across diverse clinically investigated and approved drugs, these datasets can lack specificity with respect to therapeutic indications, patient populations, and clinical contexts. As a result, Madrigal ’s safety predictions are broadly applicable but may not align with the nuanced differences across specific indications, nor fully capture clinically relevant concerns. For example, the MASH candidate firsocostat is a substrate of OATP1B1/1B3, which are transporters known to decline in patients with cirrhosis [120]. This decline could lead to differing adverse effects when firsocostat is combined with another potent OATP1B1/1B3 substrate, especially in patients at various stages of cirrhosis. This limitation positions Madrigal as a tool best suited for broad safety screening across diverse drug combinations rather than definitive, indication-specific safety evaluations. Nevertheless, Madrigal allows experts to craft queries for more focused comparisons. For example, a user could compare a new drug combination with a benchmark therapy with a known safety profile, tailored to a target population or indication. By framing such ”apples-to-apples” questions, the model can predict the relative safety of the new combination in terms of specific adverse event types. Future work could enhance this capability by integrating indication-specific information and further refining predictions for specific therapeutic contexts. In addition, factors such as the severity of the disease and prior lines of therapy should be considered.
Another limitation of Madrigal is that drug doses are not explicitly considered, which means that safety predictions are tied to the approved doses for each drug since these are the doses represented in FAERs and clinical databases with drug label information. Another fruitful avenue of future work includes incorporating clinical patient data to contextualize clinical safety models, as shown in our analysis, which incorporates patient genomic profiles. Drug administration practices, such as variable dosing schedules, drug administration sequence, and drug holidays, can be modeled to better reflect clinical practice, as personalized dosage adjustments have been linked to improved safety in clinical trials [121]. Predictions may be less reliable when a drug class is underrepresented in the training distribution, even if substantial preclinical data exist for an individual drug. For example, although adavosertib has cell viability and transcriptomic perturbation data, it is the only WEE1 inhibitor in the dataset. As a result, the model does not encounter similar perturbation patterns during training, leading to predictions that may not fully capture class-specific mechanisms, such as hematological adverse events associated with WEE1 inhibition [122, 70]. To address this, foundation models with knowledge-grounded retrieval [123, 124] that integrate preclinical literature can improve the representation of drug classes during model training.
@@
Madrigal integrates translational pharmacology with multimodal AI to predict drug combination outcomes. It identifies interactions between approved and investigational drugs, guiding co-administration to mitigate risks. In oncology and metabolic disorders, Madrigal links molecular toxicity signatures to clinical outcomes, improving precision in combination therapy selection. Madrigal paves the way for more precise drug combination design, enhancing safety assessment and therapy selection in preclinical studies and clinical research.
@@
Acknowledgements. We obtained the latest BeatAML dataset from Jeffrey Tyner, and we thank Christopher Eide for his help with data processing. We thank Man Qing Liang and Xiang Zhang for their helpful discussions on the TWOSIDES dataset and Payal Chandak for discussions on the PrimeKG knowledge graph. We thank Philip Isola, Shanghua Gao, Huan He, Wenxian Shi, and Michelle M. Li for their feedback on model development. We thank Ruth Johnson for her feedback on the paper and for discussions on using Madrigal for personalized clinical outcome prediction. We also appreciate Intae Moon’s and Walker Rickord’s comments on the manuscript. We thank Nigel Greene, Hebatallah Mohamed, and Michaël Ughetto for helpful feedback on Madrigal and analyses. We thank Jane Knöchel for interpreting model predictions in MASH. We also thank Diansong Zhou and Karthick Vishwanathan for their valuable insights into the analysis of cancer drug combinations. We gratefully acknowledge the support of NIH R01-HD108794, NSF CAREER 2339524, US DoD FA8702-15-D-0001, Harvard Data Science Initiative, Amazon Faculty Research, Google Research Scholar Program, AstraZeneca Research, Roche Alliance with Distinguished Scientists, Sanofi iDEA-iTECH, Pfizer Research, Gates Foundation (INV-079038), Chan Zuckerberg Initiative, John and Virginia Kaneb Fellowship at Harvard Medical School, Biswas Computational Biology Initiative in partnership with the Milken Institute, Harvard Medical School Dean’s Innovation Fund for the Use of Artificial Intelligence, and Kempner Institute for the Study of Natural and Artificial Intelligence at Harvard University. Any opinions, findings, conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the funders.
@@
Data availability. Data and results of our analyses are shared via the project website at https://zitniklab.hms.harvard.edu/projects/Madrigal. Datasets are also available at Harvard Dataverse repository at https://doi.org/10.7910/DVN/ZFTW3J.
@@
Authors contributions. Y.H. developed and implemented Madrigal, and designed evaluation setup. V.U. implemented first-stage pretraining of Madrigal for three modalities and integrated Madrigal with a large language model. Y.H. performed detailed analyses of Madrigal’s algorithm. X.S., V.U., and Y.H. implemented alternative methods for benchmarking. Y.H. retrieved and processed multimodal drug datasets used to train Madrigal models and prepared datasets for model evaluation, including the BeatAML, PDXE, cancer combination therapies, and MASH combination therapies. N.H. retrieved and processed cell viability data used to train Madrigal models. Y.H., X.S., and I.L. performed analyses on cancer combination therapies. I.L. processed datasets on T2D and cancer combination therapies, and Y.H. and I.L. performed analyses on polypharmacy management. L.C., D.O., B.J., and M.G. provided expertise on clinical pharmacology and safety of drug combinations. M.Z. designed and led the study. All authors contributed to writing the manuscript.
@@
Figure 1: Madrigal integrates multimodal preclinical data to predict clinical outcomes of drug combinations. a, Overview of data modalities and the modality alignment framework. The modality-specific encoders are aligned via contrastive pretraining. Madrigal extracts information from multimodal data using specialized encoders (Methods Sec. 2). b, Comprising modality-specific encoders, a fusion module, and a decoder, Madrigal is trained on expert-curated and patient-reported drug combination datasets to predict safety and efficacy outcomes. A customized transformer encoder with bottleneck tokens enhances fusion (Methods and Supplementary Fig. S1). The model enables three key applications: prediction of safety outcomes in patients receiving multiple medications, rational design of combination therapies in oncology by modeling interactions between targeted therapies and other anticancer agents, and flexible outcome description by incorporating diverse biomedical signals beyond predefined ontologies. c, The missing data modality problem is evident in the scarcity of drugs with more than two available modalities, motivating our three-stage optimization procedure, including a modality alignment step. We treat cell lines in transcriptomics as separate modalities in the fusion module (Supplementary Fig. S1), throughout the text and illustrations, we will refer to transcriptomics as a single modality to enhance readability and clarity. This distinction is intended to be self-evident from context. d, Illustration of modality-specific latent representations of ten randomly sampled drugs, shown before and after modality alignment in Madrigal. Prior to alignment, representations cluster by data type, while post-alignment, they reorganize based on drug identity, enabling cross-modal integration.
@@
Figure 2: Benchmarking Madrigal and performance analyses. a, Data splitting strategy for predicting safety outcomes of drug combinations. The split-by-drug pairs and split-by-drugs setups are illustrated for four drugs $d_{1},d_{2},d_{3}$, and $d_{4}$ with varying modality availability, and five outcomes $s_{1},s_{2},s_{3}$, $s_{4}$, and $s_{5}$. In the split-by-drugs setup, we display the training and evaluation setups. During training, all available modalities are used (for $d_{2},d_{3},d_{4}$); however, at testing, other modalities are masked (patterned boxes) for test drugs ($d_{1}$), allowing only the structure modality to be available. b, Test performance of Madrigal in the DrugBank (expert-curated drug combination outcomes) and TWOSIDES (patient-reported drug combination outcomes) datasets. In the split-by-drugs (target) dataset split, test drugs are chosen such that they have minimal overlap in their target profile with the train drugs (Methods). “w/o CL” refers to the ablation model without modality alignment; “struc. only” refers to the ablation model with only structure modality during finetuning (but with all modalities during modality alignment); and “struc. only w/o CL” refers to the ablation model without modality alignment and with only structure modality available during finetuning. AUROC: area under the receiver operating characteristic curve; AUPRC: area under the precision-recall curve; Fmax: maximum of F-measure. c,d, Test performances increase for test drugs of increasing similarity levels to train drugs in terms of structure (c) and target profile (d). The progression from “struc. only w/o CL”, “w/o CL”, to Madrigal represents progressive additions of multimodal input and modality alignment upon a simple model only taking molecular structure input. For each test drug, its structural similarity (with the train set) is calculated as the average of the highest 5 Tanimoto similarities between its Morgan Fingerprint with any train drug’s fingerprint. Target profile similarity is similarly defined as the average of the highest 5 Jaccard similarities between the drug’s target profile with any train drug’s profile. Target profiles of drugs are set of targets annotated to the drugs, obtained from DrugBank [46]. Errorbars show 95% confidence interval. Two-sided Mann-Whitney U test; **$p$-value $<0.005$. e, Test performance of Madrigal ablation with only structure modality, ablation without modality alignment (but with multi-modality), and full model, stratified by safety outcomes. Analyses in (c-e) are performed with models trained on the DrugBank dataset, split-by-drugs (random) setting. Two-sided Wilcoxon signed rank test; **$p$-value $<0.005$.
@@
Figure 3: Evaluating Madrigal predictions on external patient safety datasets. a-c, Model predictions of individual drug’s organ-specific adverse effects correlate well with concern levels in three organ-specific adverse effect datasets (drug-induced liver injury, drug-induced cardiotoxicity, drug-induced QT prolongation). Predictions of adverse effects of individual drugs are produced by duplicating the same drug as a drug pair to input to the model, and the normalized rank of each drug is then generated by ranking all such duplicate drug pairs according to their predicted scores. Error bars show 95% confidence interval. Two-sided Mann-Whitney U test; **$p$-value $<0.005$. By default, samples in neighboring concern levels are tested. No connecting lines between neighboring concern levels indicates NS. d, Model predictions of potential transporter-mediated DDIs (Methods Sec. 5.2) correlate with newly-reported transporter-mediated DDIs [57], involving doxycycline (Dox) plus multiple other drugs. Piracetam is included as a reference as it is struturally highly similar to levetiracetam. For each drug pair, the max normalized rank is the maximal normalized rank among all potential transporter-mediated DDIs among safety outcomes from DrugBank. Percentiles compare the max normalized rank of Dox + X among either X + any curated DrugBank drug (“drug + X group”) or Dox + any curated DrugBank drug (“Dox + X group”). e, Drugs sharing the same transporters, carriers, or enzymes are predicted to have a higher tendency to have relevant safety outcomes (Methods Sec. 5.2). Transporter, carrier, and enzyme information of drugs are obtained from DrugBank [46]. The highest normalized rank among all potential transporter-, carrier-, or enzyme-mediated safety outcomes is considered for each drug pair (see Methods Sec. 5.2). Two-sided Mann-Whitney U test; **$p$-value $<0.005$. f, Drugs sharing specific transporters are predicted to have a higher tendency of both common and specific transporter-related safety outcomes. Transporters recommended by regulatory authorities to examine organ-specific safety outcomes are taken from [58]. Safety outcomes shown are ranked in the highest 10 for at least one transporter (across all drug pairs sharing it). The color of each cell reflects the aggregated normalized rank (median across all drug pairs sharing corresponding transporter), and the number in each cell is the ranking (max=158) of the aggregated normalized rank of the corresponding safety outcome among all safety outcomes for drug pairs sharing the corresponding transporter. The safety profiles of drugs sharing three enzymes, carriers, and targets, respectively, are also shown for comparison.
Figure 4: (Caption on the following page.)
Figure 4. Madrigal retrieves and ranks drug combinations based on predicted safety profiles for cancer therapy. a, Distribution of predicted hematological safety risks for clinically established PARP inhibitor (PARPi) combinations. The figure compares the safety ranks of an approved combination with an acceptable safety profile (olaparib + abiraterone), a combination associated with unacceptable dose-limiting toxicities (olaparib + paclitaxel), and an investigational combination with incomplete clinical safety characterization (olaparib + ceralasertib). The normalized rank of each combination for a given safety outcome is derived by ranking the predicted probability of the outcome relative to all pairwise drug combinations. The ranking score is computed as the median normalized rank across neutropenia-related adverse outcomes (“neutropenia, increase”, “cytopenia, increase”, “myelosuppression, anemia”, “severe leukopenia, increase”, and “neutropenia, thrombocytopenia, increase”), or all hematological toxicities. The safety rank of each combination is then derived by ranking the score relative to all pairwise combinations of PARPi with another oncology agent (Methods Sec. 5.3), where a numerically lower rank indicates a more favorable safety profile. b, Predicted organ-specific toxicity profiles of PARPi-containing combinations. Organ-specific safety scores for each PARPi-drug pair (holding a PARPi fixed while varying the paired drug) are computed as the average of the three highest predicted normalized ranks for the organ system. The median of these scores across all tested drug combinations involving a specific PARPi summarizes its predicted safety profile across different organ systems. c, Comparative safety assessment across different classes of drug combinations. Left to right bars within each group represent: (1) FDA-approved drug combinations, (2) FDA-approved combinations where at least one drug is an oncology agent, (3) clinically investigated or approved PARPi combinations, (4) all tested pairwise combinations of a PARPi with another oncology agent, and (5) clinically investigated drug combinations where at least one drug is an oncology agent. Each drug combination’s safety profile is represented by the average of the five highest normalized toxicity outcome ranks for each organ system. Higher ranks indicate greater predicted safety concerns. d, Predicted safety profiles of PARPi-containing drug combinations undergoing clinical investigation or approved for clinical use. Drug combinations are ranked in ascending order based on the average of the five highest normalized toxicity outcome ranks. The (olaparib + paclitaxel) combination, which has been associated with significant hematological toxicities in clinical trials, serves as a reference for a combination with known safety concerns.

### v2

@@ Discussion
Combination therapies are central to treating complex diseases such as hypertension, cancer, and infectious diseases, yet current models often fall short in translating preclinical data to predict clinical outcomes. Integrating diverse modalities (structure, pathways, cell viability, and transcriptomics) addresses this gap. We develop Madrigal, a multimodal AI model that predicts the effects of drug combinations across 953 clinical outcomes and 21,842 compounds, including approved and investigational drugs. Madrigal surpasses single-modality and multimodal baselines for adverse drug interaction prediction, captures clinical transporter-mediated DDIs, and reflects the safety of clinically tested combinations and combinations used in chronic conditions such as T2D and MASH. Madrigal also supports personalized combination selection by predicting individualized outcomes using BeatAML genomic profiles, patient-derived xenografts, and real-world patient data.
@@
Madrigal can be applied preclinically. When prospectively scoring drug-combination risks under evaluation in ComboMATCH [89], Madrigal flags risk for eight combinations (Supplementary Fig. S8c, Supplementary Note S2). We observe correlations between proteomic changes after drug perturbation [91] and predicted DDIs (Supplementary Fig. S8a, Supplementary Note S1). These proteomic signals partially explain similarities between Madrigal drug embeddings, even after controlling for target-profile similarity (Supplementary Fig. S8b), suggesting that Madrigal captures off-target and pathway-level effects.
Multimodal models such as Madrigal can help prioritize candidates for combination testing while remaining flexible in how drugs and outcomes are encoded. This flexibility is important given limitations in drug-combination effect datasets, including incomplete annotations for outcomes that are reported clinically [92, 93] but absent from drug-effect datasets. Clinical practice patterns can also introduce biases, for example, SGLT2 inhibitors more frequently prescribed in T2D with heart failure [94], leading to different combination exposures across populations with distinct baseline risks. By linking Madrigal to an LLM interface, users can formulate free-text clinical effects that are not fully represented in standard terminologies and benchmark candidate combinations against such descriptors, potentially improving triage before experimental testing (Supplementary Fig. S9, Supplementary Note S4).
One limitation is the indication-agnostic nature of training data and predictions. Although the underlying datasets cover diverse drugs and outcomes, they often lack specificity on indication, population, and context. As a result, Madrigal ’s safety predictions are broadly applicable screens rather than indication-definitive assessments. For instance, firsocostat is an OATP1B1/1B3 substrate, and these transporter activities can decline with cirrhosis [95], potentially altering risk when co-administered with other potent OATP1B1/1B3 substrates. In practice, we envision Madrigal enabling focused comparisons that benchmark a new combination against a standard regimen with a known safety profile for a target population and incorporate indication-specific context.
Clinical failures of drug development arise primarily from insufficient efficacy (40-50%), followed by toxicity (20-30%) and pharmacokinetic issues (10-15%) [96, 97]. A favorable Madrigal safety screen is necessary but not sufficient to de-risk a combination. Decision-making can be strengthened by coupling Madrigal with efficacy models and PK/PD simulators that connect exposure to target engagement. Incorporating indication-specific efficacy datasets could support a more holistic risk-benefit predictor. Explicit dose modeling and harmonized PK parameters (currently implicit in FAERS, drug labels, and trial reports) would improve accuracy and, together with physiologically based PK modeling, extend applicability to new formulations and dose-escalation studies. Emerging efforts to compile PK parameters for combination therapies will facilitate this direction.
Another direction is deeper clinical contextualization. Incorporating richer patient data (demographics, comorbidities, concomitant medications, and genomic profiles) can refine safety prediction, as suggested by our EHR and oncology cohort analyses. Including more diverse and fine-grained clinical covariates will further improve predictive accuracy.
Madrigal presently treats safety outcomes as single end points rather than time-to-event processes, and available labels do not distinguish acute, delayed, or cumulative toxicities. Consequently, late-onset myelosuppression, chronic organ toxicity, or efficacy decay during starts/stops and dose adjustments are not modeled. Future work will leverage longitudinal cohorts and time-to-event modeling to extend predictions to dynamic, real-world treatment courses, building on our longitudinal event-time analyses. With such data, variable dosing schedules, administration sequences, and drug holidays can be modeled to better reflect clinical practice, where personalized dose adjustments have improved safety [98].
Predictions may be less reliable for underrepresented drug classes even when a specific drug has rich preclinical data. For example, adavosertib has cell-viability and transcriptomic data, but it is the only WEE1 inhibitor in the dataset; the model thus sees fewer class-consistent perturbation patterns, which may limit detection of WEE1-linked hematologic toxicities [99, 100]. Knowledge-grounded retrieval within foundation-model frameworks [101, 102, 103] could mitigate such class sparsity by integrating targeted preclinical literature during training.
@@
Madrigal integrates translational pharmacology with multimodal AI to predict drug-combination outcomes. It identifies interactions between approved and investigational agents and can guide safer co-administration. In oncology and metabolic disorders, Madrigal links molecular toxicity signals to clinical outcomes, supporting more precise selection of combination regimens. Madrigal provides a generalizable safety-screening layer that can prioritize combinations for experimental validation and inform clinical study design.
@@
Acknowledgements. We obtained the latest BeatAML dataset from Jeffrey Tyner, and we thank Christopher Eide for his help with data processing. We thank Man Qing Liang and Xiang Zhang for their helpful discussions on the TWOSIDES dataset and Payal Chandak for discussions on the PrimeKG knowledge graph. We thank Philip Isola, Shanghua Gao, Huan He, Wenxian Shi, and Michelle M. Li for their feedback on model development. We thank Ayush Noori for discussions on using Madrigal for personalized clinical outcome prediction. We appreciate Walker Rickord’s comments on the manuscript. We thank Nigel Greene, Hebatallah Mohamed, and Michaël Ughetto for helpful feedback on Madrigal and analyses. We thank Jane Knöchel for interpreting model predictions in MASH. We also thank Diansong Zhou and Karthick Vishwanathan for their valuable insights into the analysis of cancer drug combinations. We gratefully acknowledge the support of NIH R01-HD108794, NSF CAREER 2339524, US DoD FA8702-15-D-0001, Harvard Data Science Initiative, Amazon Faculty Research, Google Research Scholar Program, AstraZeneca Research, Roche Alliance with Distinguished Scientists, Sanofi iDEA-iTECH, Pfizer Research, Gates Foundation (INV-079038), Chan Zuckerberg Initiative, John and Virginia Kaneb Fellowship at Harvard Medical School, Biswas Computational Biology Initiative in partnership with the Milken Institute, Harvard Medical School Dean’s Innovation Fund for the Use of Artificial Intelligence, and Kempner Institute for the Study of Natural and Artificial Intelligence at Harvard University. Any opinions, findings, conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the funders.
@@
Data availability. Data and results of our analyses are shared via the project website at https://zitniklab.hms.harvard.edu/projects/Madrigal. Datasets are available at Harvard Dataverse repository at https://doi.org/10.7910/DVN/ZFTW3J. The BeatAML dataset is not publicly available due to patient privacy. Clinical dataset from the Dana-Farber Cancer Institute is not publicly available due to patient privacy. The process of accessing the clinical dataset from Stanford Medicine is described at https://stanford.redivis.com/datasets/53gc-8rhx41kgt. The PDXE data can be accessed via https://www.nature.com/articles/nm.3954.
@@
Authors contributions. Y.H. developed and implemented Madrigal and designed evaluation setup. Y.H. retrieved and processed multimodal drug datasets used to train Madrigal models and performed detailed analyses of Madrigal’s algorithm and model evaluation. V.U. implemented first-stage pretraining of Madrigal for three modalities and integrated Madrigal with a large language model. X.S., V.U., and Y.H. implemented alternative methods for benchmarking. N.H. retrieved and processed cell viability data. Y.H., X.S., and I.L. performed analyses on cancer combination therapies. I.L. processed datasets on T2D and cancer combination therapies, and Y.H. and I.L. performed analyses on metabolic disorders. I.M. retrieved and processed the single-index oncology cohort data. X.S. retrieved and processed the longitudinal event-time cohort data. Y.H., X.S., and I.M. performed analyses to predict personalized clinical outcomes. L.C., D.O., B.J., R.J., M.G., and A.G. provided expertise on clinical pharmacology and safety of drug combinations. L.C., D.O., and B.J. provided additional expertise on the development of new machine learning methods, design of evaluation setup, benchmarking analyses, as well as real-world applications. M.Z. designed and led the study. All authors contributed to writing the manuscript.
@@
Figure 1: Madrigal integrates multimodal preclinical data to predict clinical outcomes of drug combinations. a, Overview of data modalities and the modality alignment framework. The modality-specific encoders are aligned via contrastive learning. Madrigal extracts information from multimodal data using specialized encoders (Methods Sec. 2). b, Comprising of modality-specific encoders, a fusion module, and a prediction module, Madrigal is trained on expert-curated and patient-reported drug combination datasets to predict clinical outcomes. Attention bottleneck modules enhance fusion (Methods and Supplementary Fig. S1). The model enables three key applications: prediction of safety outcomes in patients receiving multiple medications, efficacy prediction in ex vivo studies, and personalized drug combination outcome predictions in patients. c, The missing data modality problem is evident in the scarcity of drugs with more than two available modalities. While cell lines in transcriptomics are treated as separate modalities in the fusion module (Supplementary Fig. S1), throughout the text and illustrations, we refer to transcriptomics as a single modality to enhance readability and clarity. This distinction is intended to be self-evident from context. d, UMAP of modality-specific latent embeddings of ten randomly sampled drugs, before and after modality alignment in Madrigal. Prior to alignment, embeddings cluster by data type, while post-alignment, they cluster based on drug identity, enabling cross-modal integration.
@@
Figure 2: Benchmarking Madrigal and performance analyses. a, Data splitting strategy for predicting safety outcomes of drug combinations. In the split-by-drugs setup, during training, all available modalities are used (for $d_{2},d_{3},d_{4}$); while at testing, other modalities are masked (patterned boxes) for test drugs ($d_{1}$), leaving only the structure modality available. b, Test performance of Madrigal in the DrugBank (expert-curated drug combination outcomes) and TWOSIDES (patient-reported drug combination outcomes) datasets, split-by-drugs (target) split. “W/o CL” refers to the ablation model without modality alignment; “Struc. only” refers to the ablation model with only structure modality during finetuning (but with all modalities during modality alignment); “Struc. only w/o CL” refers to the ablation model without modality alignment and with only structure modality available during finetuning. AUROC, area under the receiver operating characteristic curve; AUPRC, area under the precision-recall curve; Fmax, maximum of F-measure. c,d, Test performance increase for test drugs with increasing similarity to train drugs in terms of structure (c) or target profile (d). The progression from “Struc. only w/o CL”, “W/o CL”, to Madrigal represents progressive additions of multimodal input and modality alignment upon a simple model only taking molecular structure as input. For each test drug, its structural similarity (with the train set) is calculated as the average of the highest 5 Tanimoto similarities between its Morgan fingerprint with any train drug’s fingerprint. Target profile similarity is similarly defined as the average of the highest 5 Jaccard similarities between the drug’s target profile with any train drug’s profile. Target profiles of drugs are set of targets annotated to the drugs in DrugBank [47]. Analyses in (c-e) are performed with models trained on the DrugBank dataset, split-by-drugs (random) setting. Error bars show 95% confidence interval. Two-sided Mann-Whitney U test; **$p$-value <0.005<0.005. e, Test performance of Madrigal ablation with only structure modality, ablation without modality alignment (but with multi-modality), and full model, stratified by safety outcomes. Two-sided Wilcoxon signed rank test; **$p$-value <0.005<0.005.
@@
Figure 3: Evaluating Madrigal predictions on external patient safety datasets. a-c, Model predictions of individual drug’s organ-specific adverse effects correlate with concern levels in three organ-specific adverse effect datasets (drug-induced liver injury (a), drug-induced cardiotoxicity (b), drug-induced QT prolongation (c)). Error bars show 95% confidence interval. Two-sided Mann-Whitney U test; **$p$-value <0.005<0.005. Higher normalized rank indicates greater predicted concern. d, Model predictions of transporter-mediated DDIs (Methods Sec. 4.2) in combinations involving doxycycline (Dox). Piracetam is included as a reference as it is struturally highly similar to levetiracetam. For each drug pair, the “Normalized rank” column denotes the maximal normalized rank among all potential transporter-mediated DDIs among safety outcomes from DrugBank. Percentiles compare the max normalized rank of Dox + X among either X + any curated DrugBank drug (“drug + X group”) or Dox + any curated DrugBank drug (“Dox + X group”). e, Drugs sharing the same transporters, carriers, or enzymes are predicted to have a higher tendency to have relevant safety outcomes (Methods Sec. 4.2). Transporter, carrier, and enzyme information of drugs are obtained from DrugBank [47]. The highest normalized rank among all potential transporter-, carrier-, or enzyme-mediated safety outcomes is considered for each drug pair (Methods Sec. 4.2). Two-sided Mann-Whitney U test; **$p$-value <0.005<0.005. f, Drugs sharing specific transporters are predicted to have a higher tendency of both common and specific transporter-related safety outcomes. Safety outcomes shown are ranked in the highest 10 for at least one transporter (across all drug pairs sharing it). The color gradient reflects the aggregated normalized rank (median across all drug pairs sharing corresponding transporter), and the number in each cell is the ranking (max=158) of the aggregated normalized rank of the corresponding safety outcome among all safety outcomes for drug pairs sharing the corresponding transporter. The safety profiles of drugs sharing three enzymes, carriers, and targets, respectively, are also shown for comparison.
Figure 4: Madrigal predicts safety for clinically tested drug combinations. a, A candidate drug pair is evaluated in two ways. Top: head-to-head clinical trial arms yielding observed AE incidence. Bottom: Madrigal predicts safety scores with regard to the same adverse outcome. The model predictions are not calibrated to match the percentage. Agreement is assessed by whether the safer arm in the trial also receives a lower Madrigal score. b, Comparing Madrigal predictions with clinical trials (CT) AE data in advanced-stage clinical trials with multiple combination arms for neutropenia, hypoglycemia, anemia, and alopecia. 19/35 trials have significantly different incidences between arms for at least one of neutropenia, hypoglycemia, anemia, and alopecia. c, Comparative safety assessment across different classes of drug combinations. Left to right bars within each group represent: (1) drug combinations containing at least one cancer drug that have been investigated in advanced stage (above Phase I), (2) drug combinations indicated for cancer that have been investigated in advanced stage, (3) FDA-approved drug combinations, (4) FDA-approved drug combinations containing at least one cancer drug, (5) PARPi combinations that have been investigated in advanced stage, and (6) pairwise combinations of a PARPi with any other cancer drug. Each drug combination’s safety profile is represented by the average of the five highest normalized toxicity outcome ranks for each organ system. Higher ranks indicate greater predicted safety concerns.